Rachel invests $4,200, what is the rate of return needed to double her investment in 12 years
Found 2 solutions by jorel1380, MathTherapy:
Answer by jorel1380(3719)

The equation for this would be FV=P(1+r)^12, where FV equals future value, P is the initial principal, and r is the interest rate. So:
8400=4200*(1+r)^12
2=(1+r)^12
ln 2=ln (1+r)^12=12 ln (1+r)
0.05776226504666210911810267678818=ln (1+r)
e^0.05776226504666210911810267678818=e^ln (1+r)=(1+r)
1+r=1.0594630943592952645618252949463
r=0.0594630943592952645618252949463 or 5.9463%
☺☺☺☺

Answer by MathTherapy(10557)


Rachel invests $4,200, what is the rate of return needed to double her investment in 12 years
Depends on the number of ANNUAL compounding periods.